# Repetitive Penrose Process in Accelerating Kerr Black Holes

### v1

@@ Abstract
This paper investigates the repetitive Penrose process in accelerating Kerr black holes and explores the influence of the acceleration factor on the repetitive Penrose process. After a brief review of accelerating Kerr black holes, we study the fundamental equations of the Penrose process in this spacetime, examine the stopping conditions required for the repetitive Penrose process, and obtain corresponding numerical results. The conclusions indicate that, apart from the third law of thermodynamics similar to previous cases, accelerating Kerr black holes exhibit stronger energy extraction capabilities compared to Kerr black holes during the repetitive Penrose process. Moreover, in prior studies, the energy utilization efficiency was difficult to exceed $50\%$. However, in accelerating Kerr black holes, when the decay radius is relatively small, the energy utilization efficiency can exceed $50\%$, indicating that the reduced extractable energy primarily transforms into extracted energy rather than irreducible mass. On the other hand, when the initial value of the acceleration factor is large, the extractable energy can decrease to nearly zero, which also differs from the case of Kerr black holes in previous studies.
@@ I Introduction
The Penrose process is a mechanism for extracting energy from a rotating black hole 8 . In this process, a particle falls from infinity and decays within the ergosphere. One of the resulting particles carries negative energy and ultimately falls into the event horizon. Due to energy conservation, the other particle escapes to infinity carrying more energy than the original particle, which is equivalent to extracting rotational energy from the black hole. Since its proposal, this process has garnered widespread attention from physicists 28 ; 29 ; 30 ; 31 ; 32 ; 33 ; 34 ; 35 ; 36 . Research indicates that for energy extraction to be feasible, the velocity of the outgoing particle should exceed half the speed of light 9 ; 10 .
@@
Penrose’s pioneering work has inspired physicists to explore various alternative mechanisms for extracting energy from black holes 25 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 26 ; 18 . Recently, Ruffini et al. 19 achieved energy extraction by imposing a turning point condition on the particle’s trajectory within the original equations of motion for the Penrose process in Kerr black holes. Furthermore, they proposed a repetitive Penrose process 3 (for related early work, see 27 ). Research found that in the repetitive Penrose process, it is impossible to extract all the extractable energy from a Kerr black hole, the change in extractable energy primarily transforms into irreducible mass. This is because after each Penrose process, the new black hole mass and spin must be used for subsequent energy extraction, while also accounting for the new irreducible mass. The repetitive Penrose process is nonlinear, and the irreducible mass increases nonlinearly as well. The repetitive Penrose process proposed by Ruffini has since been extended to the repetitive electric Penrose process 6 and to Kerr-de Sitter black holes 4 . In the repetitive electric Penrose process, similarly, the total electric energy of a Reissner-Nordström black hole cannot be fully depleted, indicating that the repetitive Penrose process follows a pattern analogous to the third law of thermodynamics. In Kerr-de Sitter black holes, besides exhibiting phenomena similar to previous cases, the energy extraction capability in certain scenarios is stronger than that of Kerr black holes.
@@
The C-metric is a special class of black hole solutions in general relativity, which can describe accelerating black holes 20 and belongs to the well-known Plebański–Demiański spacetime 21 . Accelerating black holes typically refer to black holes moving with acceleration along a certain direction under specific conditions, such as being influenced by extreme gravitational sources like cosmic strings. The inclusion of an acceleration factor fundamentally alters the causal structure and asymptotic properties of the black hole spacetime. Currently, many properties of accelerating black holes have been extensively studied 22 ; 23 ; 24 ; 1 ; 2 . In this paper, we aim to further investigate the repetitive Penrose process in accelerating Kerr black holes and explore the influence of the acceleration factor on this process. In particular, we conduct a detailed study of several iterative stopping conditions. The results indicate that, apart from exhibiting a thermodynamic third law similar to previous cases, accelerating Kerr black holes possess stronger energy extraction capabilities compared to Kerr black holes during the repetitive Penrose process. Moreover, in prior research, the energy utilization efficiency was difficult to exceed $50\%$. However, in accelerating Kerr black holes, when the decay radius is relatively low, the energy utilization efficiency can exceed $50\%$, indicating that the reduced extractable energy primarily transforms into extracted energy rather than irreducible mass. On the other hand, when the initial value of the acceleration factor is large, the extractable energy can decrease to nearly zero, which also differs from previous conclusions. These findings demonstrate that conducting the repetitive Penrose process in accelerating Kerr black holes is easier and yields more extracted energy.
@@ II The Penrose Process in Accelerating Kerr Black Holes
Among these, $r_{A}$ is the acceleration horizon, while $r_{+}$ and $r_{-}$ represent the event horizon and the Cauchy horizon, respectively. In general, we have $r_{-}\leq r_{+}<r_{A}$. In this paper, to maximize energy extraction, we adopt a simplifying assumption by considering the repetitive Penrose process for an extremal accelerating Kerr black hole on the equatorial plane, where the initial black hole spin is $a=M$. This simplification has been widely employed in references 3 ; 4 .
@@ IV Repetitive Penrose Process in Accelerating Kerr Black Holes
From Table 1, it can be seen that a small portion of the reduction in extractable energy flows into the extracted energy, while the majority flows into the irreducible mass. Based on the energy utilization efficiency, it is indicated that $29.5\%$ of the change in extractable energy is converted into extracted energy, while $70.5\%$ is converted into irreducible mass. The results show that reducing the black hole’s spin cannot extract all the corresponding rotational energy. This limitation arises from the nonlinear increase in irreducible mass, which also demonstrates that the repetitive Penrose process follows a pattern analogous to the third law of thermodynamics. Furthermore, after the iteration terminates in the repetitive Penrose process, a remaining extractable energy of $0.195441M$ persists, indicating that a significant amount of energy remains to be extracted by other means. These results exhibit similarities with the case of Kerr black holes 3 .
@@ V Conclusion
Similar to previous conclusions, reducing the black hole’s spin cannot extract all the corresponding rotational energy. This limitation arises from the nonlinear increase in irreducible mass, indicating that the repetitive Penrose process follows a pattern analogous to the third law of thermodynamics. The differences lie in the fact that, at the same decay radius, as the initial $\hat{A}$ increases, the values of the energy return on investment, the energy utilization efficiency, and the extracted energy almost all increase. This demonstrates that, in the repetitive Penrose process, accelerating Kerr black holes possess stronger energy extraction capabilities compared to Kerr black holes. When the decay radius is relatively low, the energy utilization efficiency of accelerating Kerr black holes can exceed $50\%$, indicating that the reduction in extractable energy is primarily converted into extracted energy rather than irreducible mass. Furthermore, when the initial value of the acceleration factor is large, the extractable energy can decrease to nearly zero. These phenomena are entirely different from those observed in Kerr black holes.

### v2

@@ Abstract
This paper investigates the repetitive Penrose process in accelerating Kerr black holes and explores the influence of the acceleration factor on the repetitive Penrose process. After a brief review of accelerating Kerr black holes, we study the fundamental equations of the Penrose process in this spacetime, examine the stopping conditions required for the repetitive Penrose process, and obtain corresponding numerical results. The conclusions indicate that, in addition to exhibiting previously observed similar phenomena, accelerating Kerr black holes exhibit stronger energy extraction capabilities compared to Kerr black holes during the repetitive Penrose process. Moreover, in prior studies, the energy utilization efficiency was difficult to exceed $50\%$. However, in accelerating Kerr black holes, when the decay radius is relatively small, the energy utilization efficiency can exceed $50\%$, indicating that the reduced extractable energy primarily transforms into extracted energy rather than irreducible mass. On the other hand, when the initial value of the acceleration factor is large, the extractable energy can decrease to nearly zero, which also differs from the case of Kerr black holes in previous studies.
@@ I Introduction
The Penrose process is a mechanism for extracting energy from a rotating black hole 8 . In this process, a particle falls from infinity and decays within the ergosphere. One of the resulting particles carries negative energy and ultimately falls into the event horizon. Due to energy conservation, the other particle escapes to infinity carrying more energy than the original particle, which is equivalent to extracting rotational energy from the black hole. Since its proposal, this process has garnered widespread attention from physicists 28 ; 29 ; 30 ; 31 ; 32 ; 33 ; 34 ; 35 ; 36 ; 37 . Research indicates that for energy extraction to be feasible, the velocity of the outgoing particle should exceed half the speed of light 9 ; 10 .
@@
Penrose’s pioneering work has inspired physicists to explore various alternative mechanisms for extracting energy from black holes 25 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 26 ; 18 . Recently, Ruffini et al. 19 achieved energy extraction by imposing a turning point condition on the particle’s trajectory within the original equations of motion for the Penrose process in Kerr black holes. Furthermore, they proposed a repetitive Penrose process 3 (for related early work, see 27 ). Research found that in the repetitive Penrose process, it is impossible to extract all the extractable energy from a Kerr black hole, the change in extractable energy primarily transforms into irreducible mass. This is because after each Penrose process, the new black hole mass and spin must be used for subsequent energy extraction, while also accounting for the new irreducible mass. The repetitive Penrose process is nonlinear, and the irreducible mass increases nonlinearly as well. The repetitive Penrose process proposed by Ruffini has since been extended to the repetitive electric Penrose process 6 and to Kerr-de Sitter black holes 4 . In the repetitive electric Penrose process, similarly, the total electric energy of a Reissner-Nordström black hole cannot be fully depleted. In Kerr-de Sitter black holes, besides exhibiting phenomena similar to previous cases, the energy extraction capability in certain scenarios is stronger than that of Kerr black holes.
@@
The C-metric is a special class of black hole solutions in general relativity, which can describe accelerating black holes 20 and belongs to the well-known Plebański–Demiański spacetime 21 . Accelerating black holes typically refer to black holes moving with acceleration along a certain direction under specific conditions, such as being influenced by extreme gravitational sources like cosmic strings. The inclusion of an acceleration factor fundamentally alters the causal structure and asymptotic properties of the black hole spacetime. Currently, many properties of accelerating black holes have been extensively studied 22 ; 23 ; 24 ; 1 ; 2 . In this paper, we aim to further investigate the repetitive Penrose process in accelerating Kerr black holes and explore the influence of the acceleration factor on this process. In particular, we conduct a detailed study of several iterative stopping conditions. The results indicate that, in addition to the analogous characteristics previously noted, accelerating Kerr black holes possess stronger energy extraction capabilities compared to Kerr black holes during the repetitive Penrose process. Moreover, in prior research, the energy utilization efficiency was difficult to exceed $50\%$. However, in accelerating Kerr black holes, when the decay radius is relatively low, the energy utilization efficiency can exceed $50\%$, indicating that the reduced extractable energy primarily transforms into extracted energy rather than irreducible mass. On the other hand, when the initial value of the acceleration factor is large, the extractable energy can decrease to nearly zero, which also differs from previous conclusions. These findings demonstrate that conducting the repetitive Penrose process in accelerating Kerr black holes is easier and yields more extracted energy.
@@ II The Penrose Process in Accelerating Kerr Black Holes
Among these, $r_{A}$ is the acceleration horizon, while $r_{+}$ and $r_{-}$ represent the event horizon and the Cauchy horizon, respectively. In general, we have $r_{-}\leq r_{+}<r_{A}$. In this paper, to maximize energy extraction, we adopt a simplifying assumption by considering the repetitive Penrose process for an extremal accelerating Kerr black hole on the equatorial plane, where the initial black hole spin is $a=M$. Likewise, only the initial black hole is extremal; it does not remain extremal after the iterations. Indeed, it is impossible for the black hole to remain extremal throughout the repetitive Penrose process. This simplification has been widely employed in references 3 ; 4 .
@@ IV Repetitive Penrose Process in Accelerating Kerr Black Holes
From Table 1, it can be seen that a small portion of the reduction in extractable energy flows into the extracted energy, while the majority flows into the irreducible mass. Based on the energy utilization efficiency, it is indicated that $29.5\%$ of the change in extractable energy is converted into extracted energy, while $70.5\%$ is converted into irreducible mass. The results show that reducing the black hole’s spin cannot extract all the corresponding rotational energy. This limitation arises from the nonlinear increase in irreducible mass. Furthermore, after the iteration terminates in the repetitive Penrose process, a remaining extractable energy of $0.195441M$ persists, indicating that a significant amount of energy remains to be extracted by other means. These results exhibit similarities with the case of Kerr black holes 3 .
@@ V Conclusion
Similar to previous conclusions, reducing the black hole’s spin cannot extract all the corresponding rotational energy. This limitation arises from the nonlinear increase in irreducible mass. The differences lie in the fact that, at the same decay radius, as the initial $\hat{A}$ increases, the values of the energy return on investment, the energy utilization efficiency, and the extracted energy almost all increase. This demonstrates that, in the repetitive Penrose process, accelerating Kerr black holes possess stronger energy extraction capabilities compared to Kerr black holes. When the decay radius is relatively low, the energy utilization efficiency of accelerating Kerr black holes can exceed $50\%$, indicating that the reduction in extractable energy is primarily converted into extracted energy rather than irreducible mass. Furthermore, when the initial value of the acceleration factor is large, the extractable energy can decrease to nearly zero. These phenomena are entirely different from those observed in Kerr black holes.